A man has $10,000 to invest.He invests $4,000 at 5 %and $3,500 at 4%. To have a yearly income of $500 from the investment , at what rate must he invest the remainder of the money ?

Grade:10

2 Answers

MTAMBIRAJ
9 years ago
Remainder money is invested at the rate of 20%
bharat bajaj IIT Delhi
9 years ago
Say the rate of interest of remaining Rs. 2500 is x%. He will get 4000 X 5/100 from first investment and 3500 X 4/100 from second in one year, and 2500x/100 from third.
Hence, Total money from investment = 200 + 140 + 25x
340 + 25x = 500
25x = 160
x = 160/25 = 6.4 %
Thanks
Bharat Bajaj
askiitians faculty
IIT Delhi